# 2Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 3Universität Heidelberg, Interdiszipliäres Zentrum für Wissenschaftliches Rechnen, D-69120 Heidelberg, Germany 4Dipartimento di Fisica e Astronomia Galileo Galilei, Università di Padova, Vicolo dell’Osservatorio 3, I–35122 Padova, Italy

### v1

@@ 3.3.3 BBH mergers
Figures 6 and 7 show the primary and remnant mass distributions for BBHs that merge within a Hubble time. In the S23 model, no BBH merger has a primary mass $m_{1}>50\,{\rm M}_{\odot}$ for $Z>0.002$. In the C15 case, IMBH mergers can occur up to a metallicity of $Z=0.004$. In general, the BBH merger mass distribution displays a mild dependence on metallicity, with a peak at $\sim 10\,{\rm M}_{\odot}$.
The maximum merger remnant produced by these BBHs is usually two times as massive as the maximum primary BH mass. In presence of optically thick winds (model S23), the formation of IMBHs from BBH mergers is thus possible up to $Z=0.002$, where $m_{\mathrm{BH,max}}\sim 240\,{\rm M}_{\odot}$. At higher metallicities, the maximum BH remnant mass progressively decreases down to $35\,{\rm M}_{\odot}$ at $Z=0.02$.
In the C15 models, the masses delivered by BBH mergers are larger, with the possibility of having IMBH remnants up to a metallicity of $Z=0.006$.
@@ 3.3.4 Progenitors of GW231123 and GW190521
We qualitatively compare our models with the most massive BBH mergers detected by the LIGO-Virgo-KAGRA interferometers, namely GW231123 (The LIGO Scientific Collaboration et al. 2025c) and GW190521 (Abbott et al. 2020, 2020). According to our S23 model, the progenitors of GW231123 should have had a metallicity $Z\leq 0.002$. The primary mass of GW190521 (Abbott et al. 2020, 2020), instead, lies well within the mass gap predicted by both our C15 and S23 models. In a follow-up work, we will explore the dynamical formation of such mergers with our new S23 models (Torniamenti et al., in prep.).
@@ 4 Conclusions
We have investigated the impact of a new optically thick wind model (S23) on the formation of IMBHs in isolated single and binary very massive stars (VMSs). The new model is supported by several observational constraints, such as the pair-instability supernova rate (Simonato et al. 2025) and the existence of single Wolf-Rayet stars at low metallicity (Boco et al. 2025). We have integrated the evolution of VMSs with mesa (Paxton et al. 2019) and have folded our new mesa tracks inside our binary population synthesis code sevn (Iorio et al. 2023).
We find that IMBH production is suppressed at intermediate metallicity ($Z\in{}[10^{-3},0.01]$), compared to less aggressive wind models. According to our models, IMBHs cannot form from VMS collapse at the metallicity of the LMC ($Z\sim{0.008}$) and possibly even the SMC ($Z\sim{0.0025}$).
IMBHs might be able to form via stellar collisions at metallicity as high as $Z=0.004$, but only under the very optimistic assumption that each collision triggers no mass loss. However, even with this assumption for stellar collisions, the occurrence fraction of IMBHs over the total number of BHs ($\lesssim 0.2\%$) is two orders of magnitude lower compared to models with less aggressive winds.
@@
According to our S23 models, the progenitors of the most massive BBH observed by the LIGO-Virgo-KAGRA collaboration, GW231123 (The LIGO Scientific Collaboration et al. 2025c) must have had a metallicity $Z\leq 0.002$. In a follow-up study, we will explore the impact of star cluster dynamics on this result.
Figure 6: Primary BH mass distribution in BBH mergers with $M_{\mathrm{ZAMS,1}}>50{\rm M}_{\odot}$, for the S23 (upper) and C15 (lower) wind prescriptions, and for different metallicities. The blue dashed (black dotted) line displays the inferred value for GW190521 (GW231123), with the estimated uncertainty (shaded area).
@@ Acknowledgements.
We thank Gautham Sabhahit and collaborators for sharing their mesa inlists (https://github.com/Apophis-1/VMS_Paper1, https://github.com/Apophis-1/VMS_Paper2). ST acknowledges financial support from the Alexander von Humboldt Foundation for the Humboldt Research Fellowship. ST, MM and GI acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. ST, MM and LB also acknowledge financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. GI acknowledges financial support from the La Caixa Foundation for the La Caixa Junior Leader fellowship 2024. GI also acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.4, - Call for tender No. 3138 of 18/12/2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1597-1 FUGG and INST 35/1503-1 FUGG. We use the mesa software (https://docs.mesastar.org/en/latest/) version r12115; (Paxton et al. 2011, 2013, 2015, 2018, 2019). We use sevn (https://gitlab.com/sevncodes/sevn) to generate our BBHs catalogs (Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023), trackcruncher (https://gitlab.com/sevncodes/trackcruncher) (Iorio et al. 2023) to produce the tables for the interpolation. This research made use of NumPy (Harris et al. 2020), Pandas (The pandas development Team 2024), SciPy (Virtanen et al. 2020) and Matplotlib (Hunter 2007).

### v2

@@ 3.3.3 BBH mergers
Figures 6 and 7 show the primary and remnant mass distributions for BBHs that merge within a Hubble time. In the Sabhahit2023 model, no BBH merger has a primary mass $m_{1}>50\,{\rm M}_{\odot}$ for $Z>0.002$. In the Chen2015 case, IMBH mergers can occur up to a metallicity of $Z=0.004$. In general, the BBH merger mass distribution displays a mild dependence on metallicity, with a peak at $\sim 10\,{\rm M}_{\odot}$.
The maximum merger remnant produced by these BBHs is usually two times as massive as the maximum primary BH mass. In presence of optically thick winds (model Sabhahit2023), the formation of IMBHs from BBH mergers is thus possible up to $Z=0.002$, where $m_{\mathrm{BH,max}}\sim 240\,{\rm M}_{\odot}$. At higher metallicities, the maximum BH remnant mass progressively decreases down to $35\,{\rm M}_{\odot}$ at $Z=0.02$.
In the Chen2015 models, the masses delivered by BBH mergers are larger, with the possibility of having IMBH remnants up to a metallicity of $Z=0.006$.
@@ 3.3.4 Progenitors of GW231123 and GW190521
We qualitatively compare our models with the most massive BBH mergers detected by the LIGO-Virgo-KAGRA interferometers, namely GW231123 (gw231123) and GW190521 (abbottGW190521; abbottGW190521astro). According to our Sabhahit2023 model, the progenitors of GW231123 should have had a metallicity $Z\leq 0.002$. The primary mass of GW190521 (abbottGW190521; abbottGW190521astro), instead, lies well within the mass gap predicted by both our Chen2015 and Sabhahit2023 models. In a follow-up work, we will explore the dynamical formation of such mergers with our new Sabhahit2023 models (Torniamenti et al., in prep.).
4 Discussion
4.1 Impact of rotation and overshooting
In this work, we investigate the impact of a recent model accounting for optically thick winds (Sabhahit2023) on the formation of IMBHs from VMSs. We have considered non-rotating stellar models with a fixed overshooting parameter. However, both stellar rotation and core overshooting may have a relevant impact on the final BH mass. In a follow-up paper, we will include these additional parameters in detail. In the following, we briefly discuss their expected impact on our results.
Stellar rotation affects both the envelope and the core mass, leading to enhanced stellar-mass loss during the main sequence (langer1998, but see also muller2014), and triggering chemical mixing in the core (Sabhahit2023). In the optically thin regime, rotation generally results in enhanced mass loss when the critical rotation is approached ($\Omega/\Omega_{\mathrm{c}}>0.4$). As a consequence, it can further quench the BH production at low metallicity $Z<0.01\,Z_{\mathrm{\odot}}$, where optically thick winds are not activated (winch2024; winch2025). Also, rotation can drive the transition to the optically-thick regime at lower initial masses (Sabhahit2023; boco2025). For VMS, this mainly affects the evolution at low $Z$, where the star can retain a large fraction of its angular momentum due to weaker mass loss. At high $Z$, wind mass loss can efficiently remove the angular momentum, leading to a minor impact from stellar rotation on the final BH mass.
Core overshooting also triggers chemical mixing, affecting the final core mass similarly to rotation (Sabhahit2022). In the optically-thin regime, increasing the overshooting parameter leads to more massive cores, with larger values $M_{\mathrm{BH}}$ for the same $M_{\mathrm{ZAMS}}$ (winch2024). For VMSs, optically thick winds already lead to the formation of a fully mixed star if the initial stellar mass is large enough ($M_{\mathrm{ZAMS}}>200\,{\rm M}_{\odot}$, Sabhahit2022), while removing the stellar envelope. As a consequence, overshooting is not expected to play an important role on the final BH mass.
In summary, we expect overshooting and rotation to affect the relation between $M_{\mathrm{ZAMS}}$ and $M_{\mathrm{BH}}$ in the low-metallicity and low-mass end of the parameter space considered, where mass loss is driven by optically-thin winds. In such regimes, a larger overshooting parameter and a fast rotation may further reduce the maximum value of $Z$ at which IMBHs are produced. When more vigorous stellar winds are present, VMS evolution is expected to mainly be driven by stellar mass loss, with a less relevant impact from these parameters.
4.2 Caveats of VMS population models
In our population-synthesis models, we have simply drawn primary stellar masses from a Kroupa IMF between 50 and 500 M, corresponding to the extreme high-mass end of the mass function. In most star forming regions, the presence of stars in this mass range is dominated by stochastic effects or possibly by additional quenching effects (e.g., Weidner2006). Thus, the conclusions of our population-synthesis simulations only apply to extreme star-forming regions, such as the R136 cluster in the LMC (Crowther2010; bestenlehner2020), the super-star clusters in starburst galaxies (e.g., the Antennae galaxies, Whitmore1995) and the high-redshift massive clusters recently discovered by the James Webb Space Telescope (Vanzella2023; Adamo2024; Mowla2024). Moreover, it is also possible that VMSs are the result of multiple stellar collisions in very dense star clusters (Portegies2002). In such case, population studies require dedicated NN-body simulations (e.g., dicarlo2021; Rantala2024).
@@ 5 Conclusions
We have investigated the impact of a new optically thick wind model (Sabhahit2023) on the formation of IMBHs in non-rotating isolated single and binary very massive stars (VMSs). The new model is supported by several observational constraints, such as the pair-instability supernova rate (simonato2025) and the existence of single Wolf-Rayet stars at low metallicity (boco2025). We have integrated the evolution of VMSs with mesa (Paxton2019) and have folded our new mesa tracks inside our binary population synthesis code sevn (Iorio2023).
We find that IMBH production is suppressed at intermediate metallicity ($Z\in{}[10^{-3},0.01]$), compared to less effective wind models. According to our models, IMBHs cannot form from VMS collapse at the metallicity of the LMC ($Z\sim{0.008}$) and possibly even the SMC ($Z\sim{0.0025}$).
IMBHs might be able to form via stellar collisions at metallicity as high as $Z=0.004$, but only under the very optimistic assumption that each collision triggers no mass loss. However, even with this assumption for stellar collisions, the occurrence fraction of IMBHs over the total number of BHs ($\lesssim 0.2\%$) is two orders of magnitude lower compared to models with less effective wind mass-loss.
@@
According to our Sabhahit2023 models, the progenitors of the most massive BBH observed by the LIGO-Virgo-KAGRA collaboration, GW231123 (gw231123) must have had a metallicity $Z\leq 0.002$. In a follow-up study, we will explore the impact of star cluster dynamics on this result.
Figure 6: Primary BH mass distribution in BBH mergers with $M_{\mathrm{ZAMS,1}}>50{\rm M}_{\odot}$, for the Sabhahit2023 (upper) and Chen2015 (lower) wind prescriptions, and for different metallicities. The blue dashed (black dotted) line displays the inferred value for GW190521 (GW231123), with the estimated uncertainty (shaded area).
@@ Acknowledgements.
We thank the anonymous referee for their insightful comments and suggestions. We thank Gautham Sabhahit and collaborators for sharing their mesa inlists (https://github.com/Apophis-1/VMS_Paper1, https://github.com/Apophis-1/VMS_Paper2). We thank Guglielmo Costa and Kendall Shepherd for useful comments and discussions. ST acknowledges financial support from the Alexander von Humboldt Foundation for the Humboldt Research Fellowship. ST, $M$ and GI acknowledge financial support from the European Research Council for the ERC Consolidator grant DEMOBLACK, under contract no. 770017. ST, $M$ and LB also acknowledge financial support from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) STRUCTURES. GI acknowledges financial support from the La Caixa Foundation for the La Caixa Junior Leader fellowship 2024. GI also acknowledges financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.4, - Call for tender No. 3138 of 18/12/2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. EK and $M$ acknowledge support from the PRIN grant METE under contract No. 2020KB33TP. EK acknowledges financial support from the Fondazioni Gini for the Gini grant. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grants INST 35/1597-1 FUGG and INST 35/1503-1 FUGG. We use the mesa software (https://docs.mesastar.org/en/latest/) version r12115; (Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019). We use sevn (https://gitlab.com/sevncodes/sevn) to generate our BBHs catalogs (Spera2019; Mapelli2020; Iorio2023), trackcruncher (https://gitlab.com/sevncodes/trackcruncher) (Iorio2023) to produce the tables for the interpolation. The parsec stellar tracks that we used for this work are publicly available at http://stev.oapd.inaf.it/PARSEC/ (costa2025). This research made use of NumPy (Harris20), Pandas (Pandas2024), SciPy (SciPy2020) and Matplotlib (Hunter2007).